# Scientific Literature Hierarchy Construction and Organization

### v3

@@ Abstract
We motivate Science Hierarchography, the goal of organizing scientific literature into a high-quality hierarchical structure that spans multiple levels of abstraction—from broad domains to specific studies. Such a representation can provide insights into which fields are well-explored and which are under-explored. To achieve this goal, we develop a hybrid approach that combines efficient embedding-based clustering with LLM-based prompting, striking a balance between scalability and semantic precision. Compared to LLM-heavy methods like iterative tree construction, our approach achieves superior quality-speed trade-offs. Our hierarchies capture different dimensions of research contributions, reflecting the interdisciplinary and multifaceted nature of modern science. We evaluate its utility by measuring how effectively an LLM-based agent can navigate the hierarchy to locate target papers. Results show that our method improves interpretability and offers an alternative pathway for exploring scientific literature beyond traditional search methods. 111 Code, data and demo are available: https://github.com/JHU-CLSP/science-hierarchography
@@
Muhan Gao, Jash Shah, Weiqi Wang, Daniel Khashabi Department of Computer Science Johns Hopkins University Baltimore, MD 21218, USA {mgao38,jshah48,wwang194,danielk}@jhu.edu
@@ 2 Related Work
Gierarchy induction: The field of taxonomy induction has progressed from early pattern-based techniques to modern LLM-augmented methods. Seminal work by Hearst (1992) introduced the use of hand-crafted hyponym patterns for extracting is-a relationships. Subsequent research expanded on this using statistical methods and large-scale information extraction to identify hypernym-hyponym structures (Pantel and Pennacchiotti, 2006; Yang and Callan, 2009; Girju et al., 2006).
@@
Our work differs in scope, scale, and methodological design. We focus on scaling taxonomy induction for the domain of scholarly literature—a setting that presents greater challenges than typical setups (e.g., entity hierarchy) due to the complexity, size, and evolving nature of scientific content. The most comparable effort is by Oarga et al. (2024), though our broader objectives require fundamentally different algorithmic strategies and operate without access to ground truth labels.
@@
In Table 1 we summarize the differences with existing hierarchical resources. While most prior systems are limited to one/two level(s) of depth and rely on manually assigned labels for indexing—a process often prone to bias (Hadfield, 2020)—our approach supports deeper, algorithmically generated hierarchies with semantically rich node descriptions. This enables a more flexible and interpretable representation of scientific knowledge.
@@ 3 Science Hierarchography: Toward Hierarchy of Scholarly Work
We begin with a formal problem definition (§3.1), followed by content representation (§3.2) and depth considerations (§3.3).
@@ 3.1 Formal Problem Statement
We define the task of Science Hierarchography as an inference problem where the input is a large set of scientific papers: $P=\{p_{1},p_{2},\ldots,p_{n}\}$. The goal is to infer a hierarchical structure for these papers (e.g., their problems and techniques). This structure represents levels of specificity and abstraction, with nodes closer to the root representing broader topics. Broader topics are at the upper levels, while more specific subtopics and individual papers are at the lower levels.
Formally, this goal is defined as inferring a graph that encodes relationships between nodes (atomic concepts representing scholarly ideas or goals). This is formally defined as $T\vcentcolon=(N,E)$ where:
  • NN is the set of all concepts (nodes) present in the hierarchy: $N=\{n_{1},n_{2},\ldots,n_{N}\}$.
  • $E$ is the set of all edges connecting node pairs: $E=\{r_{1},r_{2},\ldots,e_{M}\}$ where each edge is a node pair $r_{i}=(n_{p},n_{c})$ connects a parent node $n_{p}$ to its child node $n_{c}$.
  • For each edge $e\in E$, there is a type attribute: $\text{type}(e)\in\{\text{``isA''},\text{``instanceOf''}\}$ associated with the edge.
  • Each node $n\in N$ has a property $n.prop$ which defines the content of the node (e.g., a sentence description or a few keywords).
The edge (relations connecting two nodes) have types that is either “isA” relationships or “instanceOf.” The "isA" relationship defines a hierarchical link between node pairs, indicating a child node is a subclass of its more abstract parent node (e.g., "RLHF isA RL" means "RLHF" is a type of "RL"). The "instanceOf" relationship indicates that a specific instance belongs to a concept category (e.g., "Ouyang et al. (2022) instanceOf RLHF" means "Ouyang et al. (2022)" is a specific instance of the broader “RLHF” cluster).
@@ 3.2 Decomposing Papers to Contributions
We use an LLM to preprocess each paper (title and abstract) and break them down into a pre-defined set of contributions, akin to prior work Hope et al. (2017); Chan et al. (2018) that mine “problem schema” from existing documents. We consider the following contribution types: (1) problem statement (the problem addressed), (2) solution (the technical approach used), (3) result (the key finding), and (4) topic (the overarching themes). (See §C for prompts and examples). We note that each contribution may include additional dimensions (sub-contributions). For instance, a “result” encompasses both the “outcome” and its “potential impact.”
@@ 3.3 Choosing Hierarchy Depth
While the ideal number of hierarchy layers is ultimately empirical, we can build useful intuition from the structure of a near-balanced tree. For a tree with branching factor $b$ and depth $L$, the total number of nodes is roughly $O(b^{L})$. To organize $C$ contributions, the number of nodes should scale with $C$, implying a depth of $L=O(\log_{b}C)$. In practice, we use $L=3$ for a 2K-paper corpus and $L=4$ for 10K papers, consistent with this logarithmic scaling. Extrapolating further, corpora of $10^{7}$ papers would likely require depths of $L=6$ or $7$.
@@ 4 Tackling Science Hierarchography
We present algorithms to address our proposed goal. We start with our main method, Scychic4.1), explore its special cases (§4.2), and then describe alternative baselines that rely more heavily on LLMs (fLMSci; §4.3). While all approaches leverage LLMs to some extent, they differ significantly in their reliance on them: some require many calls (linear or quadratic in the number of papers), while others are more efficient (e.g., logarithmic). Since our goal is to scale to millions of papers, minimizing LLM usage is critical. Our objective is to identify the method that yields the highest-quality hierarchy with the lowest LLM overhead, balancing quality, latency, and cost.
4.1 Scychic: Alternating Between Clustering and Summarization
This approach is based on the following design choices: (1) access to embedder, a neural model that converts a description into a $d$-dimensional vector, (ideally) capturing its semantic meaning; (2) a clustering algorithm clusterer that, given the hyperparameter $k$, generates $k$ clusters; (3) a contribution type (e.g., problem definition) and its dimensions $C$ extracted per paper as detailed in §3.2 which determines the focus of the node descriptions; (4) summarizer, an LLM that generates a summary description which (ideally) provides a more abstract description of a collection of node descriptions; and (5) the total number of hierarchy layers $L$ and target number of clusters in each layer $(k_{1},k_{2},\ldots,k_{L})$.
Initialization: The approach begins by embedding each papers. For each paper $p_{i}$, we embed each component in $C$: $\texttt{{\color[rgb]{0.09765625,0.0078125,0.6796875}embedder}}(c_{j}^{i})\in\mathbb{R}^{d}$, where $j\in C$. This process results in $|C|$ embeddings per paper. We concatenate these embeddings, yielding $\mathbb{R}^{d.|C|}$ embeddings per paper. With this, we present the main algorithm which involves two phases:
Top-down phase: First, we use a top-down strategy, iterating through the top half of the layers. For each layer $l\in(1\ldots\left\lfloor L/2\right\rfloor)$ (starting from the root and moving downward), we cluster the documents using clusterer.
For the first layer ($L=1$), we partition the entire paper collection into $k_{1}$ top-level clusters using their embeddings. For the second level, we process each top-level cluster independently, i.e., within each top-level cluster, we take only the papers belonging to that cluster and apply clusterer to their embeddings to divide them into subgroups.
This process continues recursively: for subsequent layers, clustering operates on the data points within each cluster. We run as many clustering algorithms as there are clusters in the previous layer. The allocation of the total cluster count across different top-level clusters is determined proportionally based on the number of papers in each top-level cluster. If a top-level cluster contains a larger proportion of papers, it will be allocated more subclusters from the total cluster count. This proportional allocation ensures that areas with higher paper density receive finer-grained clustering.
This recursively subdivides each cluster at level $l$ into multiple clusters at level $l+1$. This top-down phase continues until we reach level $\left\lfloor L/2\right\rfloor$. Alongside clustering, we also use the summarizer to generate abstracted summaries for each of the clusters. The generated cluster description follows the same structure or style as the input descriptions. For example, if the inputs are statements about problem categories, the output from summarizer is also in the same style, but more abstract.
Bottom-up phase: In the second phase, we switch to a bottom-up strategy to construct the remaining levels ($\left\lfloor L/2\right\rfloor+1$ through $L$). To form clusters for bottom-level (layer $L$), we apply clusterer to the paper embeddings within each sub-cluster within level-$\left\lfloor L/2\right\rfloor$ (the lowest level clustering obtained from top-down approach). We then use the summarizer to create an abstracted description for each cluster.
We repeat this process for all layers from $L$ to $\left\lfloor L/2\right\rfloor+1$. To build layer $l$, we start by embedding the generated cluster summaries from the level below l1l-1 using embedder, similar to how we embedded the papers. We then run the clustering clusterer on these new embeddings and generate abstracted summaries for the clusters to group these summaries into higher-level clusters. This bottom-up aggregation continues until we connect with the previously constructed level $\left\lfloor L/2\right\rfloor$ clusters.
@@ 4.1 Scychic: Alternating Between Clustering and Summarization
2:Initialization: For each paper $p_{i}\in P$, embed their selected components to form $\mathbb{R}^{d\times|C|}$.
@@
10:for each cluster $\kappa$ at level $\lfloor L/2\rfloor$ do $\triangleright$ Bottom-up phase
@@
13:     Collect the embeddings of papers within $\kappa$.
@@ 4.3 Pure LLM-based Baselines
We introduce baselines that heavily utilize LLM calls, based on the hypothesis that LLMs can make high-quality local decisions, collectively forming a robust global structure. The potential cost here is the need to make many LLM calls. We refer to these baselines as fLMSci (pronounced “flimsy”) and present two variants below.
Initializing a Seed Hierarchy: The first step involves creating a seed hierarchy, starting with the hierarchy of sciences from the Wikipedia page on branches of science.222en.wikipedia.org/wiki/Branches_of_science We made several adjustments to this hierarchy, detailed in §D, where the resulting seed hierarchy is also included.
fLMSci (parallel): Adding Many Items in Parallel with Few Prompts: This approach involves adding items to the seed hierarchy in parallel. First, gather all unique items extracted from different papers and create batches of 100 items each (batching because all of these items would not fit within the context window of LLM). Then, implement a multi-threaded program where each thread adds a batch of keywords to a clone of the seed hierarchy. Finally, merge (with a Python program, not an LLM call) all these cloned trees into a single comprehensive structure.
fLMSci (incremental): Expand the tree by incrementally adding items one at a time: This approach involves an iterative, layer-by-layer prompt that navigates the tree starting from the root node and performs specific actions. By default, the supported actions include: (a) Go down: select one of the nodes to traverse down to a lower layer in the tree; (b) Add sibling: Insert a new node at the same level; (c) Make parent: Create a new parent node; (d) Discard: Ignore the item, if no suitable location is found. This prompt are shown in Fig.11.
The available actions vary depending on the tree’s current position. To prevent placing detailed items too high in the tree (above layer 3), we hide the actions that involve nodes (b) and (c). This restriction provides more control and reduces errors, as observed in pilot experiments. If the traversal reaches a child node, we hide the option to move further down (a). Another issue identified was frequent mistakes in earlier layers, likely due to broad labels where multiple categories could fit. To address this, we provide more context for the first layer by replacing label strings with descriptive definitions, as shown in Fig.10.
@@ 5.2 Evaluation as Utilization
Two key design choices guide our evaluation: (a) selecting appropriate queries, and (b) choosing a reliable LLM. For (a), we sample paper titles and abstracts as queries. Although we considered generating language questions from papers, pilot studies showed both approaches yield similar results, so we use the simpler method. For (b), we use Qwen2.5-32b-instruct, which demonstrated performance closest to GPT-4 among open models in our pilot evaluations (§B).
Starting at the top of the hierarchy, we provide the evaluator (LLM) with the query and the list of cluster descriptions (prompt in Fig. 2). The model must select the cluster it deems most relevant. We move to the next level only if the chosen cluster contains the target paper. At each subsequent level, the same query is used, but only subclusters of the previously selected (and correct) cluster are shown. This continues until the model reaches the level containing the target paper. We report two metrics: Strict Accuracy (Strict-Acc), which captures how often the model selects the exact correct node, and Layer-1 Accuracy (L1-Acc), which measures how often it selects the correct top-level subtree.

### v4

# Science Hierarchography: Hierarchical Organization of Science Literature

@@ Abstract
We motivate Science Hierarchography, the goal of organizing scientific literature into a high-quality hierarchical structure that spans multiple levels of abstraction—from broad domains to specific studies. Such a representation can provide insights into which fields are well-explored and which are under-explored. To achieve this goal, we develop a hybrid approach that combines efficient embedding-based clustering with LLM-based prompting, striking a balance between scalability and semantic precision. Compared to LLM-heavy methods like iterative tree construction, our approach achieves superior quality-speed trade-offs. Our hierarchies capture different dimensions of research contributions, reflecting the interdisciplinary and multifaceted nature of modern science. We evaluate its utility by measuring how effectively an LLM-based agent can navigate the hierarchy to locate target papers. Results show that our method improves interpretability and offers an alternative pathway for exploring scientific literature beyond traditional search methods.
@@
Muhan Gao1, Jash Shah2, Weiqi Wang2, Kuan-Hao Huang1, Daniel Khashabi2 1Texas A&M University, 2Johns Hopkins University
@@ 2 Related Work
Hierarchy induction: The field of taxonomy induction has progressed from early pattern-based techniques to modern LLM-augmented methods. Seminal work by Hearst (1992) introduced the use of hand-crafted hyponym patterns for extracting is-a relationships. Subsequent research expanded on this using statistical methods and large-scale information extraction to identify hypernym-hyponym structures (Pantel and Pennacchiotti, 2006; Yang and Callan, 2009; Girju et al., 2006).
@@
Our work differs in scope, scale, and methodological design. We focus on scaling taxonomy induction for the domain of scholarly literature—a setting that presents greater challenges than typical setups (e.g., entity hierarchy) due to the complexity, size, and evolving nature of scientific content.
The closest works are Oarga et al. (2024), which build domain-specific hierarchies (e.g., Chemical Elements) using iterative LLM refinement, and Zhu et al. (2025), which organizes survey-based collections of fewer than 100 papers. Our work differs in scope and scale: our broader objectives require fundamentally different algorithmic strategies and operate without access to ground truth labels.
@@
In Table 1 we summarize the differences with existing hierarchical resources. While most prior systems are limited to one/two level(s) of depth and rely on manually assigned labels for indexing—a process often prone to bias (Hadfield, 2020). For example, Scopus employs a fixed two-layer hierarchy (ASJC codes) assigned at the journal level, so paper-level classifications are inherited rather than content-derived. In contrast, our approach supports deeper, algorithmically generated hierarchies with semantically rich node descriptions. This enables a more flexible and interpretable representation of scientific knowledge. We provide a detailed comparison with existing systems in §A.
@@ 3 Science Hierarchography: Toward Hierarchy of Scholarly Work
We begin with the problem definition (§3.1), followed by content representation (§3.2) and depth considerations (§3.3).
@@ 3.1 Formal Problem Statement
We define the task of Science Hierarchography as an inference problem where the input is a large set of scientific papers: $P=\{p_{1},p_{2},\ldots,p_{n}\}$. The goal is to infer a hierarchical structure (i.e., a tree) for a specific contribution type (e.g., problem statement) of a collection of papers. The nodes of this tree are the atomic concepts representing scholarly ideas or goals. The edge (relations connecting two nodes) encode whether one node is a specific version of another node (i.e., “isA” relationships) which defines a hierarchical link between node pairs, indicating a child node is a subclass of its more abstract parent node (e.g., “RLHF isA RL” means “RLHF” is a type of “RL”). The specific papers $P$ are the nodes of this tree. The overall hierarchy represents levels of specificity and abstraction, with nodes closer to the root representing broader topics. Broader topics are at the upper levels, while more specific subtopics and individual papers are at the lower levels.
Why a tree structure? A tree offers a clear, interpretable way to capture hierarchical relations among scientific ideas, showing how concepts specialize or generalize without cycles or ambiguity. Each node inherits meaning from its ancestors, tracing the progression from broad themes to concrete contributions. We build one tree per contribution type (e.g., problem, method; §3.2), forming a forest of hierarchies that naturally accommodates the multi-faceted and interdisciplinary nature of scientific research.
@@ 3.2 Decomposing Papers to Contributions
We use an LLM (gpt-4o-2024-08-06) to preprocess each paper (title and abstract), use the paper’s title and abstract as the input, and request the LLM to break them down into a pre-defined set of contributions, akin to prior work Hope et al. (2017); Chan et al. (2018) that mines “problem schema” from existing documents. We consider the following contribution types: (1) problem statement (the problem addressed), (2) solution (the technical approach used), (3) result (the key finding), and (4) topic (the overarching themes). (See §C for prompts and examples). We note that each contribution may include additional dimensions (sub-contributions). For instance, a “result” encompasses both the “outcome” and its “potential impact.” In total, this yields $C=11$ sub-contributions per paper. The LLM performs consistently during extraction: when we deliberately remove information from the input (primarily from the abstract), it correctly leaves the corresponding sub-contributions blank rather than hallucinating content, which demonstrates its reliability.
@@ 3.3 Choosing Hierarchy Depth
While the ideal number of hierarchy layers is ultimately empirical, we can build useful intuition from the structure of a near-balanced tree. For a tree with branching factor $b$ and depth $L$, the total number of nodes is roughly $O(b^{L})$. To organize $C$ contributions, the number of nodes should scale with $C$, implying a depth of $L=O(\log_{b}C)$. In practice, we use $L=3$ for a 2K-paper corpus and $L=4$ for 10K papers, consistent with this logarithmic scaling. Extrapolating further, corpora of $10^{7}$ papers would likely require depths of $L=6$ or 77.
@@ 4 Tackling Science Hierarchography
We present algorithms to address our proposed goal. We start with our main method, Scychic4.1), explore its special cases (§4.2), and then describe alternative baselines that rely more heavily on LLMs (fLMSci; §4.3). While all approaches leverage LLMs to some extent, they differ significantly in their reliance on them: some require many calls (linear or quadratic in the number of papers), while others are more efficient (e.g., logarithmic). Since our goal is to scale to a large number of papers, minimizing LLM usage is critical. Our objective is to identify a method that yields the highest-quality hierarchy with the lowest LLM overhead, balancing quality, latency, and cost.
@@
2:Initialization: For each paper $p_{i}\in P$, using embedder embed their selected components to form $\mathbb{R}^{d\times|C^{\prime}|}$.
@@
10:for each cluster $\tau$ at level $\lfloor L/2\rfloor$ do $\triangleright$ Bottom-up phase
@@
13:     Collect the embeddings of papers within $\tau$.
@@
4.1 Scychic: Alternating Between Clustering and Summarization
Overview: Our method builds each contribution-type hierarchy through two complementary stages: a top-down phase that clusters paper embeddings into progressively finer subgroups and summarizes each cluster, followed by a bottom-up phase that embeds and reclusters the generated summaries to form higher-level abstractions. The combined process yields coherent, interpretable hierarchies that capture both fine-grained and global structure.
Ingredients: This approach is based on the following design choices: (1) access to embedder, a neural model that converts a description into a $d$-dimensional vector, (ideally) capturing its semantic meaning; (2) a clustering algorithm clusterer that, given the hyperparameter $k$, generates $k$ clusters; (3) a contribution type (e.g., problem definition) and its dimensions $C^{\prime}$ extracted per paper as detailed in §3.2 which determines the focus of the node descriptions; (4) summarizer, an LLM that generates a summary description which (ideally) provides a more abstract description of a collection of node descriptions; and (5) the total number of hierarchy layers $L$ and target number of clusters in each layer $(k_{1},k_{2},\ldots,k_{L})$.
Specifically, for embedder we use gte-Qwen2-7B-instruct, for our summarizer we use Llama-3.3-70B-Instruct Grattafiori et al. (2024), and for clusterer, we apply k-means clustering. (further details in §G.)
Initialization: The approach begins by embedding each paper. For each paper $p_{i}$, we embed each component in $C^{\prime}$: $\texttt{{\color[rgb]{0.09765625,0.0078125,0.6796875}embedder}}(c_{j}^{i})\in\mathbb{R}^{d}$, where $j\in C^{\prime}$. This process results in $|C^{\prime}|$ embeddings per paper. We concatenate these embeddings, yielding $\mathbb{R}^{d.|C^{\prime}|}$ embeddings per paper. We now present the main algorithm consisting of two phases:
Phase 1: Top-down: We begin with a top-down strategy that recursively partitions the paper set through the upper half of the hierarchy ($l\in[1,\left\lfloor L/2\right\rfloor]$). At the first level, all papers are clustered into $k_{1}$ groups using their embeddings. Each cluster is then processed independently—papers within a cluster are reclustered using clusterer to form finer subgroups. The number of subclusters assigned to each parent cluster scales linearly with its paper count, ensuring denser regions of the corpus receive finer resolution. This recursive subdivision continues until level $\left\lfloor L/2\right\rfloor$, producing a coarse-to-fine hierarchy. At each level, summarizer to generate abstracted summaries for each of the clusters based on the clustered papers’ titles and abstracts. The generated cluster description follows the same structure or style as the input descriptions. For example, if the inputs are statements about problem categories, the summaries are also in the same style, but more abstract.
Phase 2: Bottom-up: In the second phase, we switch to a bottom-up strategy to construct the remaining levels ($\left\lfloor L/2\right\rfloor+1$ through $L$). To form clusters for bottom-level (layer $L$), we apply clusterer to the paper embeddings within each sub-cluster within level-$\left\lfloor L/2\right\rfloor$ (the lowest level clustering obtained from the top-down approach). We then use the summarizer to create an abstracted description for each cluster. We repeat this process for all layers from $L$ to $\left\lfloor L/2\right\rfloor+1$. To build layer $l$, we start by embedding the generated cluster summaries from the level below l1l-1 using embedder, similar to how we embedded the papers. We then run the clustering clusterer on these new embeddings and generate abstracted summaries for the clusters to group these summaries into higher-level clusters. This bottom-up aggregation continues until we connect with the previously constructed level $\left\lfloor L/2\right\rfloor$ clusters.
@@ 4.3 Pure LLM-based Baselines
We introduce baselines that heavily utilize LLM calls, based on the hypothesis that LLMs can make high-quality local decisions, collectively forming a robust global structure. The potential cost here is the need to make many LLM calls. We refer to these baselines as fLMSci (pronounced “flimsy”) and present two variants below. For both methods, we use gpt-4o-2024-08-06 to extract the contributions (§3.2), and Llama-3.3-70B-Instruct to place them into the hierarchy.
Initializing a Seed Hierarchy: The first step involves creating a seed hierarchy, starting with the hierarchy of sciences from the Wikipedia page on branches of science111en.wikipedia.org/wiki/Branches_of_science and refined through several adjustments detailed in §D.
fLMSci (parallel): parallel addition of contributions: This approach expands the seed hierarchy in parallel using a small number of LLM calls. All unique contributions extracted from papers are first collected and divided into batches of 100 (to fit within the LLM’s context window). A multi-threaded program then assigns each batch to a separate thread, where the LLM adds those contributions to a cloned copy of the seed hierarchy. Finally, the cloned hierarchies are merged (via a Python script rather than additional LLM calls) into a single unified structure.
fLMSci (incremental): Incremental tree expansion: This method builds the hierarchy iteratively by adding one contribution at a time through layer-by-layer prompting. Starting from the root, the model navigates the tree and performs one of four actions: (a) Go down: move to a lower-level node; (b) Add sibling: insert a new node at the same level; (c) Make parent: create a new parent node; or (d) Discard: ignore the contribution if no suitable location exists (Fig. 11). Available actions depend on the current position in the tree. To avoid placing detailed contributions too high in the tree, we disable node-creation actions (b, c) above layer 3. When reaching a leaf node, the Go down action (a) is also unavailable. Pilot studies revealed frequent early-layer errors due to broad category labels; to mitigate this, we replaced top-level labels with descriptive definitions (Fig. 10), improving contextual understanding and placement accuracy.
@@ 5.2 Evaluation as Utilization
Our evaluation design involves two choices: (a) queries and (b) an evaluation model. For (a), we sample paper titles and abstracts. Although we considered generating language questions from papers, pilot studies showed both approaches yield similar results, so we use the simpler method. For (b), we use Qwen2.5-32b-instruct, which performed closest to GPT-4 among open models (§B.1).
The process starts at the root: given a query and cluster descriptions (Fig. 2), the LLM selects the most relevant cluster. If it contains the target paper, traversal continues recursively through subclusters until the correct paper-level node is reached. We report two metrics: Strict-Acc, the fraction of cases where the model finds the target node, and L1-Acc, which measures how often it correctly identifies the top-level subtree containing the target.
Validation: We also validate the reliability of our LLM-based evaluation through both (a) human assessment and (b) evaluation based on existing human-annotated hierarchy from ORKGB.2).